Keith’s Florists has 16 delivery trucks, used mainly to deliver flowers and flower arrangements in the Greenville, South Carolina, area. Of these 16 trucks, 6 have brake problems. A sample of 6 trucks is randomly selected. What is the probability that 2 of those tested have defective brakes? (Round your answer to 4 decimal places.)
Answer by stanbon(75887)

Keith’s Florists has 16 delivery trucks, used mainly to deliver flowers and flower arrangements in the Greenville, South Carolina, area. Of these 16 trucks, 6 have brake problems. A sample of 6 trucks is randomly selected. What is the probability that 2 of those tested have defective brakes? (Round your answer to 4 decimal places.)
----
Binomial Problem with n = 6 and p(defect) = 6/16 = 3/8
-----------
P(x = 2) = 6C2(3/8)^2*(5/8)^4 = binompdf(6,3/8,2) = 0.3219
=========================
Cheers,
Stan H.
=============